# Mitigating Suboptimality of Deterministic Policy Gradients in Complex Q-functions

### v1

@@ 4.4 Approximate Surrogate Functions
    $\tilde{\mu}_{M}(s)$ is the action selected by the current maximizer actor $\mu_{M}$, having a high Q-value,
@@
    $\nu_{i}(s;a_{<i})$ is the action proposed by the $i$-th actor conditioned on previous actions $a_{<i}$,
$\mathcal{L}_{\text{approx}}=\mathbb{E}_{s\sim\rho^{\mu_{M}}}\left[\sum_{a\in\{\tilde{\mu}_{M}(s),\nu_{i}(s;a_{<i})\}}\left\|\hat{\Psi}_{i}(s,a;a_{<i})-\Psi_{i}(s,a;a_{<i})\right\|_{2}^{2}\right].$ (9)
This design ensures $\hat{\Psi}_{i}$ is updated on high Q-value actions and thus the landscape is biased towards those values. This makes the gradient flow trend in the direction of high Q-values. So, even when $a_{i}$ from $\nu_{i}$ falls in a region of zero gradients for $\Psi_{i}$, in $\hat{\Psi}_{i}$ would provide policy gradient in a higher Q-value direction, if it exists. Figure 5 shows $\Psi_{1}$ and $\hat{\Psi}_{1}$ in restricted inverted pendulum task.
@@ 4.5 SAVO-TD3 Algorithm and Design Choices
     Evaluate $a_{0}=\mu(s),\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}{a_{i}=\nu_{i}(s;a_{<i})}$
     Add perturbations with OU Noise $\hat{a}_{i}=a_{i}+\epsilon_{i}$
@@
     Update $Q,Q_{2}\leftarrow r+\gamma\min\{Q^{\prime}(s^{\prime},\tilde{a}),Q_{2}^{\prime}(s^{\prime},\tilde{a})\}$
     Update $\hat{\Psi}_{i}$ with Eq. 9 $\forall i=1,\dots k$
@@
     Update actor $\nu_{i}$ with Eq. 8 $\forall i=1,\dots k$
@@ 2. Exploration in Additional Actors:
Added actors $\nu_{i}$ explore the surrogate landscapes for high-reward regions by adding OU (Lillicrap et al., 2015) or Gaussian (Fujimoto et al., 2018) noise to their actions.
@@ 4. Conditioning on Previous Actions:
Actors $\nu_{i}$ and surrogates $\hat{\Psi}_{i}$ are conditioned on preceding actions via FiLM layers (Perez et al., 2018) as in Figure 4.
@@ 5. Discrete Action Space Tasks:
We apply 1-nearest-neighbor $f_{\text{NN}}$ before Q-value evaluation to ensure the Q-function is only queried at in-distribution actions.
SAVO-TD3 employs SAVO actor to systematically reduce the local optima in its base algorithm TD3. We empirically validate the proposed design improvements.
@@ 5 Environments
We evaluate SAVO on discrete and continuous action space environments with challenging Q-value landscapes. More environment details are presented in Appendix C and Figure 13.
Locomotion in Mujoco. We evaluate on MuJoCo (Todorov et al., 2012) environments of Hopper-v4, Walker2D-v4, Inverted Pendulum-v4, and Inverted Double Pendulum-v4.
@@
Locomotion in Restricted Mujoco. We create a restricted locomotion suite of the same environments as in MuJoCo. A hard Q-landscape is realized via high-dimensional discontinuities that restrict the action space. Concretely, a set of predefined hyper-spheres (as shown in Figure 6) in the action space are sampled and set to be valid actions, while the other invalid actions have a null effect if selected. The complete details can be found in Appendix C.3.1.
Adroit Dexterous Manipulation. Rajeswaran et al. (2017) propose manipulation tasks with a dexterous multi-fingered hand. Door: In this task, a robotic hand is required to open a door with a latch. The challenge lies in the precise manipulation needed to unlatch and swing open the door using the fingers. Hammer: the robotic hand must use a hammer to drive a nail into a board. This task tests the hand’s ability to grasp the hammer correctly and apply force accurately to achieve the goal. Pen: This task involves the robotic hand manipulating a pen to reach a specific goal position and rotation. The objective is to control the pen’s orientation and position using fingers, which demands fine motor skills and coordination.
Mining Expedition in Grid World. We develop a 2D Mining grid world environment (Chevalier-Boisvert et al., 2018) where the agent (Figure 13) navigates a 2D maze to reach the goal, removing mines with correct pick-axe tools to reach the goal in the shortest path. The action space includes navigation and tool-choice actions, with a procedurally-defined action representation space. The Q-landscape is non-convex because of the diverse effects of nearby action representations.
@@ 6.1 Effectiveness of SAVO in challenging Q-landscapes
We compare SAVO against the following baseline actor architectures, where $k=3$:
@@
  • 1-Actor, $k$ samples (Wolpertinger): Gaussian sampling centered on actor’s output. For discrete actions, we select $k$-NN discrete actions around the continuous action (Dulac-Arnold et al., 2015).
  • $k$-Actors (Ensemble): Best Q-value among diverse actions from ensemble (Osband et al., 2016).
  • Evolutionary actor (CEM): Iterative CEM search over the action space (Kalashnikov et al., 2018).
@@
  • SAVO: Our method with $k$ successive actors and surrogate Q-landscapes.
@@
  • SAVO - Approximation: removes the approximate surrogates (§4.4), using $\Psi_{i}$ instead of $\hat{\Psi}_{i}$.
@@
  • SAVO + Action Smoothing: TD3’s policy smoothing (Fujimoto et al., 2018) compute Q-targets.
@@
Aggregate performance. We utilize performance profiles (Agarwal et al., 2021) to aggregate results across different environments in Figure 7(a) (evaluation mechanism detailed in Appendix G.1). SAVO consistently outperforms baseline actor architectures like single-actor (TD3) and sampling-augmented actor (Wolpertinger), showing wide robustness across challenging Q-landscapes. In Figure 7(b), SAVO outperforms its ablations, validating each proposed component and design decision.
Per-environment results. In discrete action tasks, the Q-value landscape is only well-defined at exact action representations and nearby discrete actions might have very different values (§3.2.1). This makes the Q-value landscape uneven, with multiple peaks and valleys (Figure 2). For example, actions in Mining Expedition involve both navigation and tool-selection which are quite different, while RecSim and RecSim-Data have many diverse items to choose from. Methods like Wolpertinger that sample many actions a local neighborhood perform better than TD3 which considers a single action (Figure 8). However, SAVO achieves the best performance by directly simplifying the non-convex Q-landscape. In restricted locomotion, there are several good actions that are far apart. SAVO actors can search and explore widely to optimize the Q-landscape better than only nearby sampled actions. Figure 16 ablates SAVO in all 7 environments and shows that the most critical features are its successive nature, removing policy smoothing, and approximate surrogates.
@@ 6.2 Q-Landscape Analysis: Do successive surrogates reduce local optima?
Figure 9 visualizes surrogate landscapes in Inverted-Pendulum-Restricted for a given state $s$. Successive pruning and approximation smooth the Q-landscapes, reducing local optima. A single actor gets stuck at a local optimum $a_{0}$ (left), but surrogate $\hat{\Psi}_{1}$ uses $a_{0}$ as an anchor to find a better optimum $a_{1}$. The maximizer policy finally selects the highest Q-valued action among $a_{0},a_{1},a_{2}$. Figure 24 extends this visualization to Inverted-Double-Pendulum-Restricted.  Figure 23 shows how one actor is sufficient in the convex Q-landscape of unrestricted Inverted-Pendulum-v4. Figures 2526 show how Hopper-v4 Q-landscape provides a path to global optimum, while Hopper-Restricted is non-convex.
@@ 6.3 Challenging Dexterous Manipulation (Adroit)
In Adroit dexterous manipulation tasks (Door, Pen, Hammer) (Rajeswaran et al., 2017), SAVO improves the sample efficiency of TD3 (Figure 10). The non-convexity in Q-landscape likely arises from nearby actions having high variance outcomes like grasping, missing, dropping, or no impact.
Figure 10: SAVO improves the sample-efficiency of TD3 on Adroit dexterous manipulation tasks.
@@ 6.4 Quantitative Analysis: The Effect of Successive Actors and Surrogates
We investigate the effect of increasing the number of successive actor-surrogates in SAVO in Pendulum (Figure 11(a)) and MineWorld (Figure 11(b)). Additional actor-surrogates significantly help to reduce severe local optima initially. However, the improvement saturates as the suboptimality gap reduces.
(a) Length Analysis Pendulum
(b) Length analysis Mine World
(c) Return Improvement
Figure 11: (L) More successive actor-surrogates are better, (R) SAVO v/s single-actor on inference.
Next, we show that successive actors are needed because a single actor can get stuck in local optima even with an optimal Q-function. In Figure 11(c), we consider a SAVO agent trained to optimality with 3 actors. When we remove the additional actors, the remaining single-actor agent resembles TD3 trained to maximize an “optimal” Q-function. However, the significant performance gap indicates that the single actor could not find optimal actions for the given Q-function.
6.5 Does RL with Resets address the issue of Q-function Optimization?
Figure 12: Reset (primacy bias) does not improve Q-optimization.
Primacy bias (Nikishin et al., 2022; Kim et al., 2024) occurs when an agent is trapped in suboptimal behaviors from early training. To mitigate this, methods like resetting parameters and re-learning from the replay buffer aim to reduce reliance on initial samples. We run TD3 in MineEnv with either a full-reset or last-layer reset every 200k, 500k, or 1 million iterations. None of these versions outperformed the original TD3 algorithm without resets. This is because resetting does not help an actor to navigate the Q-landscape better and can even cause an otherwise optimal actor to get stuck in a suboptimal solution during retraining. In contrast, the SAVO actor architecture specifically addresses the non-convex Q-landscapes, being a more robust method to finding closer to optimal actions.
@@ 6.6 Further experiments to validate SAVO
  • Unrestricted locomotion. Figure 15 shows that both SAVO and baselines achieve optimal performance in simple Q-landscapes, confirming effective hyperparameter tuning (§G.4, §G.3) and indicating that the baselines underperform due to the complexity introduced in Q-landscapes.
  • SAVO orthogonal to SAC. Figure 17 shows that SAVO+TD3 > SAC > TD3, indicating that SAC’s stochastic policy does not address non-convexity, but can itself suffer from local optima (Figure 18)
  • Design Choices. Figure 20 shows that LSTM, DeepSet, and Transformers are all valid choices as summarizers of preceding actions $a_{<i}$ in SAVO. Figure 21 shows that FiLM conditioning on $a_{<i}$ especially helps for discrete action space tasks but has a smaller effect in continuous action space. In Figure 22(a), we find Ornstein-Uhlenbeck (OU) noise and Gaussian noise to be largely equivalent.
  • Massive Discrete Actions. SAVO outperforms in RecSim with 100k and 500k actions (Figure 19).
@@
Table 1: Compute v/s Performance Gain (Mujoco)
@@ 7 Limitations and Conclusion
Introducing more actors in SAVO has negligible influence on GPU memory, but leads to longer inference time (Table 1). However, even for 3 actor-surrogates, SAVO achieves significant improvements in all our experiments. Further, for tasks with a simple convex Q-landscape, a single actor does not get stuck in local optima, making the gain from SAVO negligible. In conclusion, we improve Q-landscape optimization in deterministic policy gradient RL with Successive Actors for Value Optimization (SAVO) in both continuous and large discrete action spaces. We demonstrate with quantitative and qualitative analyses how the improved optimization of Q-landscape with SAVO leads to better sample efficiency and performance.
Acknowledgements
This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant (No.RS-2019-II190075, Artificial Intelligence Graduate School Program, KAIST) and National Research Foundation of Korea (NRF) grant (NRF-2021H1D3A2A03103683, Brain Pool Research Program), funded by the Korea government (MSIT). Ayush Jain was supported partly as intern at Naver AI Lab during the initiation of the project. We appreciate the fruitful discussions with members of CLVR Lab at KAIST and LiraLab at USC.
Appendix
[…]
Appendix A Proof of Convergence of Maximizer Actor in Tabular Settings
Theorem A.1 (Convergence of Policy Iteration with Maximizer Actor).
In a finite Markov Decision Process (MDP) with finite state space $\mathcal{S}$, consider a modified policy iteration algorithm where, at each iteration $n$, we have a set of $k+1$ policies $\{\nu_{0},\nu_{1},\dots,\nu_{k}\}$, with $\nu_{0}=\mu_{n}$ being the policy at the current iteration learned with DPG. We define the maximizer actor $\mu_{M}$ as:
$\mu_{M}(s)=\arg\max_{a\in\{\nu_{0}(s),\nu_{1}(s),\dots,\nu_{k}(s)\}}Q^{\mu_{n}}(s,a),$ (10)
where $Q^{\mu_{n}}(s,a)$ is the action-value function for policy $\mu_{n}$. Then, the modified policy iteration algorithm using the maximizer actor is guaranteed to converge to a final policy $\mu_{N}$.

### v2

@@ 4.2.2 Approximately track Q-function with a bias towards high valued actions
    Tracking: $\tilde{\mu}_{M}(s)$ represents the action taken in the environment at which the latest online update to the Q-function has been made following Eq. 1, which helps track the value of $\Psi_{i}$.
@@
    High-value Bias: $\nu_{i}(s;a_{<i})$ is the action proposed by the $i$-th actor conditioned on previous actions $a_{<i}$, which is expected to be a high-valued action.
This design ensures $\hat{\Psi}_{i}$ is updated on high Q-value actions and thus the landscape is biased towards those values. This makes the gradient flow trend in the direction of high Q-values. So, even when $a_{i}$ from $\nu_{i}$ falls in a region of zero gradients for $\Psi_{i}$, in $\hat{\Psi}_{i}$ would provide policy gradient in a higher Q-value direction, if it exists. Figure 5 shows $\Psi_{1}$ and $\hat{\Psi}_{1}$ in restricted inverted pendulum task. Figure 23 analyzes $\mathcal{L}_{\text{approx}}$ over training, demonstrating that $\hat{\Psi}_{i}$ stays close to $\Psi_{i}$ while smoothing it.
4.2.3 Successive Gradient-based Actors for Each Surrogate Optimization
To effectively reduce local optima using the approximate surrogates $\hat{\Psi}_{1},\dots,\hat{\Psi}_{k}$, we design the policies $\nu_{i}$ to optimize their respective $\hat{\Psi}_{i}(s,a;\,a_{<i})$. Each $\nu_{i}$ focuses on regions where $Q(s,a)\geq\max_{j<i}Q(s,a_{j})$, allowing it to find better optima than previous policies. The actor $\nu_{i}$ is conditioned on previous actions $\{a_{0},\dots,a_{i-1}\}$, summarized via deep sets (Zaheer et al., 2017) (see Figure˜4). The maximizer actor $\mu_{M}$ (Eq. 5) then selects the best action among these proposals.
We train each actor $\nu_{i}$ using gradient ascent on its approximate surrogate $\hat{\Psi}_{i}$, similarly to DPG:
$\nabla_{\phi_{i}}J(\phi_{i})\;=\;\mathbb{E}_{s\sim\rho^{\mu_{M}}}\Bigl[\nabla_{a}\,\hat{\Psi}_{i}\bigl(s,a;\,a_{<i}\bigr)\Big|_{a=\nu_{i}(s;\,a_{<i})}\;\cdot\;\nabla_{\phi_{i}}\,\nu_{i}\bigl(s;\,a_{<i}\bigr)\Bigr].$ (9)
@@ 4.3 SAVO-TD3 Algorithm and Design Choices
  Evaluate $a_{0}=\mu(s),\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}{a_{i}=\nu_{i}(s;a_{<i})}\;\forall i=i\dots k$
  Perturb with OU Noise $\hat{a}_{i}=a_{i}+\epsilon_{i}\;\forall i=i\dots k$
@@
[…]
  Update $\hat{\Psi}_{i}$ with Eq. 8 $\forall i=1\dots k$
@@
  Update actor $\nu_{i}$ with Eq. 9 $\forall i=1\dots k$
@@ 2. Exploration in Additional Actors:
Successive actors $\nu_{i}$ explore their surrogate landscapes by adding OU (Lillicrap et al., 2015) or Gaussian (Fujimoto et al., 2018) noise to their outputs, effectively discovering high-reward regions.
@@ 4. Conditioning on Previous Actions:
Actors $\nu_{i}$ and surrogates $\hat{\Psi}_{i}$ are conditioned on preceding actions via FiLM layers (Perez et al., 2018) as in Fig. 4.
@@ 5. Discrete Action Space Tasks:
We apply 1-nearest-neighbor $f_{\text{NN}}$ before the Q-function, so it is only queried at in-distribution actions. For gradient flow into the actor, a noisy Q-function is added. See Q-smoothing in §14.6.
SAVO-TD3 systematically reduces local optima through successive surrogates while leveraging TD3 as a robust RL baseline. In the next section, we validate these design choices through experiments, demonstrating SAVO-TD3’s effectiveness in complex reinforcement learning tasks against alternate actor architectures.
@@ 5 Environments
We evaluate SAVO on discrete and continuous action space environments with challenging Q-value landscapes. More environment details are presented in Appendix 11 and Figure 12.
Locomotion in Mujoco-v4. We evaluate Mujoco (Todorov et al., 2012) environments of Hopper, Walker2D, Inverted Pendulum, and Inverted Double Pendulum without any restrictions.
Locomotion in Restricted Mujoco. We create a restricted locomotion suite of the same environments as in Mujoco-v4, which results in hard Q-landscapes due to high-dimensional discontinuities from a restricted action space. Concretely, a set of predefined hyper-spheres (as shown in Figure 6) in the action space are sampled and set to be valid actions, while the other invalid actions have a null effect if selected. The complete details of the mobility restrictions on the action space can be found in Appendix 11.3.1.
Adroit Dexterous Manipulation (Rajeswaran et al., 2017) Door: In this task, a robotic hand is required to open a door with a latch. The challenge lies in the precise manipulation needed to unlatch and swing open the door using the fingers. Hammer: the robotic hand must use a hammer to drive a nail into a board. This task tests the hand’s ability to grasp the hammer correctly and apply force accurately to achieve the goal. Pen: This task involves the robotic hand manipulating a pen to reach a specific goal position and rotation. The objective is to control the pen’s orientation and position using fingers, which demands fine motor skills and coordination.
Mining Expedition in Grid World. We develop a 2D Mining grid world environment (Chevalier-Boisvert et al., 2018) where the agent (Appendix Fig. 12) navigates a 2D maze to reach the goal, removing mines with correct pick-axe tools to reach the goal in the shortest path. The action space includes navigation and tool-choice actions, with a procedurally-defined action representation space. The Q-landscape is non-convex because of the diverse effects of nearby action representations.
@@ 6 Experiments
6.1 Effectiveness of SAVO in challenging Q-landscapes
We compare SAVO against the following baseline actor architectures:
@@ 6.1 Effectiveness of SAVO in challenging Q-landscapes
  • 1-Actor, kk=3 samples (Wolpertinger): Gaussian sampling centered on actor’s output. For discrete actions, we select $3$-NN discrete actions around the continuous action (Dulac-Arnold et al., 2015).
  • kk=3-Actors (Ensemble): Each actor (Osband et al., 2016) can find different local optima, improving the best action.
  • Evolutionary actor (CEM): Repeated rounds of search with CEM over the action space (Kalashnikov et al., 2018).
@@
  • SAVO: Our method with $3$ successive actors and surrogate Q-landscapes.
@@
  • SAVO - Approximation: removes the approximate surrogates (Sec. 4.2.2), using $\Psi_{i}$ instead of $\hat{\Psi}_{i}$.
@@
  • SAVO + Action Smoothing: TD3’s policy smoothing (Fujimoto et al., 2018) adds action noise to compute Q-targets.
@@
Aggregate performance. We utilize performance profiles (Agarwal et al., 2021) to aggregate results across different environments in Figure 7(a) (evaluation mechanism detailed in Appendix 15.1). SAVO consistently outperforms baseline actor architectures like single-actor (TD3) and sampling-augmented actor (Wolpertinger), showing the best robustness across challenging Q-landscapes. In Figure 7(b), SAVO outperforms its ablations, validating each proposed component and design decision.
Per-environment results. In Mining Expedition, the action space has semantically different navigation and tool-choice actions, while RecSim and RecSim-Data have a large and diverse set of items. The Q-landscape is significantly non-convex in such discrete tasks because the continuous action goes through a nearest-neighbor step to select a discrete item. Thus, sampling more neighbors in a local neighborhood via Wolpertinger is better than TD3’s single action in Figure 8. However, the optimal action is not necessarily near the initial guess. Therefore, SAVO achieves the best performance by directly addressing global non-convexity. In restricted locomotion with a discontinuous action space, SAVO’s actors can search far separated regions to optimize the Q-landscape better than only nearby sampled actions. Appendix Figure 22 ablates SAVO in all 7 environments and shows that the most critical features are its successive nature, removing policy smoothing, and approximate surrogates.
@@ 6.2 Q-Landscape Analysis: Do successive surrogates reduce local optima?
Figure 10: TD3 is improved with SAVO on Adroit dexterous manipulation tasks.
Figure 11: SAVO’s improvement scales well when additional actor-surrogates are added until its performance saturates and completely mitigates the suboptimality of TD3. While the gains are diminishing beyond 3-5 actors in the environments we considered, 10 actors are mostly enough to produce optimal performance (shown in red). For RecSim, which is an especially non-convex Q-landscape because of 10,000 actions and a 45-D action representation space, we note that increasing to 15 actors achieves the optimal performance.
In Figure 9, we visualize the surrogate landscapes in Inverted Pendulum-Restricted for one state $s$. Due to successive pruning and approximation, the Q-landscapes become smoother with reduced local optima. A single actor gets stuck in a severe local optimum $a_{0}$. However, surrogate $\Psi_{1}$ utilizes $a_{0}$ as an anchor and finds a better (global) optimum $a_{1}$. The maximizer policy selects $a_{0},a_{1},\text{or }a_{2}$, whichever has the highest Q-value. Appendix Figure 28 shows that convex Q-landscapes are easily optimized, while Figure 29 shows how SAVO successfully optimizes the non-convex Q-landscapes in all other tasks. Further analysis can be found in Appendix 18.2.
@@ 6.3 Challenging Dexterous Manipulation (Adroit)
In Adroit (Rajeswaran et al., 2017) dexterous manipulation on Door, Pen, and Hammer, we compared SAVO to TD3  (Fujimoto et al., 2018) and observed that SAVO successfully addressed the Q-landscape challenges in TD3 algorithm (Figure 10) and TD3 has been improved with SAVO.
@@ 6.4 Quantitative Analysis: The Effect of Successive Actors and Surrogates
We investigate the effect of increasing the number of successive actor-surrogates in SAVO in Figure 11. Additional actor-surrogates significantly help to reduce severe local optima initially. However, the improvement saturates as the suboptimality gap reduces. While we still report main SAVO results using 3 actors, SAVO significantly improves with 10 actors (Figure 11, Figure 26) across tasks.
@@ 6.5 Further experiments to validate SAVO
  • Baseline Optimization. Figure 15 shows that baselines are fairly optimized, on par with SAVO on tasks with a simple Q-landscape. Hyperparameter optimization details are discussed in §15.3.
  • SAVO orthogonal to SAC. Figure 18 shows that SAVO+TD3 > SAC > TD3; thus, SAC’s stochastic policy does not address TD3’s non-convexity. In fact, SAC also suffers from local optima (§17, Figure 27) that SAVO+SAC mitigates successfully in unrestricted Ant-v4 and Half-Cheetah-v4.
  • Design Choices. Figure 20 shows that LSTM, DeepSet, and Transformers are all valid choices as summarizers of successive actions $a_{<i}$ in SAVO. Figure 21 shows that FiLM conditioning on $a_{<i}$ helps in discrete action spaces, but affects continuous action space less. For exploration, we compared Ornstein-Uhlenbeck (OU) noise and Gaussian noise and found them to be largely equivalent across all baselines (Figure 17). In §14.7, we tried specialized initializations to enforce diversity in the SAVO’s actors and surrogates but did not observe major gains.
  • Massive Discrete Actions. SAVO also improves in RecSim-100k and RecSim-500k (Figure 19).
  • Resetting baselines. SAVO outperforms resetting techniques (Nikishin et al., 2022; Kim et al., 2024) in addressing local optima, as shown in Figure 16.
@@
Table 1: Compute v/s Performance Gain
@@ 7 Limitations and Conclusion
Introducing more actors in SAVO has negligible influence on GPU memory, but leads to longer inference time (Table 1). However, even for 3 actor-surrogates, SAVO achieves significant improvements in all our experiments. Further, for tasks with a simple convex Q-landscape, single actors do not get stuck in local optima, reducing the improvements with SAVO. In conclusion, we improve Q-landscape optimization in actor-critic RL with Successive Actors for Value Optimization (SAVO) in both continuous and large discrete action spaces. We demonstrate with quantitative and qualitative analyses how the improved optimization of Q-landscape with SAVO leads to better sample efficiency and performance.
Supplementary Materials
The following content was not necessarily subject to peer review.
8 Reproducibility
With the aim of promising the reproducibility of our results, we have attached our code in the supplementary materials, which contain all environments and all baseline methods we report in the paper. The specific commands to reproduce all baselines across all environments are available in README. We have also included all relevant hyperparameters and additional details on how we tuned each baseline method in Appendix Table 3.
9 Proof of Convergence of Maximizer Actor in Tabular Settings
Theorem 9.1 (Convergence of Policy Iteration with Maximizer Actor).
Consider a modified policy iteration algorithm where, at each iteration, we have a set of $k+1$ policies $\{\nu_{0},\nu_{1},\dots,\nu_{k}\}$, with $\nu_{0}=\mu$ being the current policy learned with DPG under its assumptions of finite states, continuous actions, and regularity conditions. We define the maximizer actor $\mu_{M}$ as:
$\mu_{M}(s)=\arg\max_{a\in\{\nu_{0}(s),\nu_{1}(s),\dots,\nu_{k}(s)\}}Q(s,a).$ (10)
In the tabular setting, the modified policy iteration algorithm using the maximizer actor converges to the locally optimal policy.